tan(theta) tan(120-theta) tan(120+theta)= 1 by root 3 then value of thetaa is

Grade:12th pass

1 Answers

Roma
6 years ago
Tan A . Tan (120- A ) . Tan (120+ A) =  1 by root 3
Tan 3A = 1 by root 3
3A = n( pie )  + ( pie ) / 6
A = ( n ( pie ) / 3 ) + ((pie ) / 18 ) = (n + 1/6 ) (pie/ 3) = ( 6n +1) (pie/ 18 )
Here ( pie ) = angle in trigonometry equal to the value of 180 degree